# PersonalView: Multi-View Consistent Human Image Customization via In-Context Learning

### v1

@@ Abstract
Recent advances in personalized generative models demonstrate impressive results in creating identity-consistent images of the same person under diverse settings. Yet, we note that most methods cannot control the viewpoint of the generated image, nor generate consistent multiple views of the person. To address this problem, we propose a lightweight adaptation method, PersonalView, capable of enabling an existing model to acquire multi-view generation capability with as few as 100 training samples. PersonalView consists of two key components: First, we design a conditioning architecture to take advantage of the in-context learning ability of the pre-trained diffusion transformer. Second, we preserve the original generative ability of the pretrained model with a new Semantic Correspondence Alignment Loss. We evaluate the multi-view consistency, text alignment, identity similarity, and visual quality of PersonalView and compare it to recent baselines with potential capability of multi-view customization. PersonalView significantly outperforms baselines trained on a large corpus of multi-view data with only 100 training samples.
Figure 1: PersonView compared to conventional customization methods. PersonalView generates personalized images consistent with multiple views given one reference image. Conventional methods like PULID (Guo et al., 2024) have limited control over the viewpoint in the prompt (i.e., left, middle, and right view) and do not have multi-view consistency.
@@ 1 Introduction
Recent years have witnessed an explosion of research endeavors in the domain of human image customization (Li et al., 2024e; Guo et al., 2024; Wang et al., 2024; Ye et al., 2023; Li et al., 2024c; 2025; Yang et al., 2024). These methodologies exhibit the capability for human customization to user-provided photographic input. For example, given a user’s personal photograph, such approaches can synthesize novel customizations depicting the individual seated on a beach or smiling in a grassland setting. Despite the recent surge in human image customization techniques, a significant challenge persists in achieving multi-view consistent customization.
Method
Geometry
Consistency
Visual
Consistency
w/o Large
MV Dataset
w/o Test-Time
Training
Human Identity
Consistency
Prompt
Control
Conventional Customization
(e.g., PuLID (Guo et al., 2024))
Image to Multi-View
(e.g., Era3D (Li et al., 2024d))
Unified Multi-Modal Model
(e.g., BAGEL (Deng et al., 2025))
CustomDiffusion360
(Kumari et al., 2024)
PersonalView
@@
What if a user wishes to change the view of the customized human while simultaneously synthesizing it in a novel context, as exemplified in Figure 1? Simply, employing viewpoint-specific prompts like ‘from the left view’ has limited control on customized images from diverse angles. Furthermore, the resulting outputs often suffer from a lack of inter-view coherence including geometry consistency of the human body and visual consistency manifesting across a range of visual attributes, encompassing body gestures, facial features, expressions, background elements, apparel, and accessories. Consequently, a salient inquiry emerges: how can multi-view consistency be preserved during the customized image generation process to satisfy user requirements?
@@
In this work, we introduce a new task: Multi-View Customization (MVC) of human images conditioned on a single user-provided photograph. Without test-time tuning, it empowers users to generate multi-view personalized images exhibiting robust geometry consistency and visual consistency across different views with promising identity fidelity and versatile prompt-based control, as shown in Table 1. It not only enables users to explicitly control viewpoints for greater flexibility in customization and visual creation, but also holds potential for extension to other critical domains such as 3D modeling and reconstruction.
@@
The core of this task is how to control the generation of multiple views while maintaining geometric and visual consistency. We are inspired by the in-context learning ability of the DiT-based models, such as FLUX (Labs, 2024b). As shown in prior work (Huang et al., 2024; Kang et al., 2025), they have the ability to generate a grid of images with roughly consistent visual content. This observation motivates our hypothesis that the model has a strong prior in multi-view capability, though the consistency is not guaranteed. Therefore, we propose to exploit this 3D-aware prior by in-context learning to foster more explicit multi-view consistency.
@@
To this end, we propose employing in-context depth-maps as the cues to activate FLUX’s capacity for geometry and visual consistent generation. Specifically, we introduce a novel generation framework based on multi-view depth-maps, termed PersonalView. At first, our goal is to obtain multi-view depth maps that are consistent with the provided prompts, which will serve as contextual cues for further processing. To achieve this, in the first stage, we leverage a pre-trained customization model (Guo et al., 2024) to perform preliminary sampling. Subsequently, we employ the SMPL model (Goel et al., 2023) for fitting and rendering, generating multi-view depth maps that align with the desired customization. Then, these multi-view human depth-maps are arranged into a four-panel grid, serving as an in-context conditioning signal for the depth-conditioned model to synthesize multi-view consistent customized images.
To enhance the geometry and visual consistency across multiple views, we further fine-tune the depth-conditioned adapter with a multi-view in-context learning paradigm. However, given the limited diversity of scenes within multi-view data, the model is susceptible to overfitting, potentially compromising its prompt controllability. To mitigate this, we introduce a Semantic Correspondence Alignment Loss specifically designed for DiT-based models to preserve the original model’s semantic control capabilities. Concretely, this loss incorporates a frozen branch of the original model during training. By aligning the correspondence between text tokens and visual tokens extracted from both the trainable and frozen branches, it effectively retains the original model’s semantic control without impeding its in-context learning abilities.
@@
Our proposed PersonalView consistently preserves multi-view human identity while demonstrating strong adherence to editing prompts, requiring only a few training steps without large-scale multi-view human dataset. It exhibits robust multi-view consistency across various attributes, including facial features, body gestures, background, and clothing, while concurrently preserving excellent prompt controllability and identity fidelity. Extensive experimentation demonstrates its superior performance compared to existing novel view synthesis methods that typically necessitate extensive training on large-scale multi-view datasets.
@@
  • We introduce a novel task, Multi-View Customization (MVC) conditioned on a single user-provided photograph without test-time tuning, which aims to generate multi-view consistent customized images with precise identity and diverse prompt-control.
  • We propose PersonalView, leveraging in-context multiview depth maps to activate the in-context learning capabilities of DiT-based models with only a few training samples, facilitating the generation of geometry and visual consistent human images.
  • We introduce a Semantic Correspondence Alignment loss tailored for DiT-based models, which preserves the original model’s semantic control capabilities without impeding its in-context learning abilities.
@@ 2.1 Image Customization
In the domain of Text-to-Image (T2I) generation, a variety of approaches have emerged to address identity (ID) customization (Gal et al., 2022; Li et al., 2024e; Gal et al., 2024; Valevski et al., 2023; Xiao et al., 2024; Ma et al., 2024; Peng et al., 2024; Li et al., 2023; 2024a). A seminal method in this line of work is Textual Inversion (Gal et al., 2022), which encodes user-specific identity information into a dedicated token embedding while keeping the T2I model parameters fixed. To improve identity fidelity, In contrast, encoder-based paradigms aim to directly inject identity representations into the generation pipeline. For instance, PhotoMaker (Li et al., 2024e) leverages large-scale identity datasets to construct robust ID embeddings from diverse image samples. Similarly, PuLID (Guo et al., 2024) introduces a more precise ID supervision mechanism by minimizing identity loss between synthesized outputs and reference images. Recently, CustomDiffusion360 enables explicit control over object viewpoints in the customization of text-to-image diffusion models; however, it requires test-time training and is limited to object-level customization rather than human subjects.
@@ 2.2 Multi-View Images Synthesis
Cross-view consistency plays a pivotal role in multi-view generation. MVFusion (Tang et al., 2023) initiates this direction by parallel multi-view image generation with correspondence-aware attention, which facilitates cross-view information exchange and supports textured scene mesh reconstruction. Building upon it, subsequent works (Tseng et al., 2023; Kant et al., 2024; Gu et al., 2024; Li et al., 2024b; Shen & Tang, 2024) incorporate epipolar constraints into diffusion models to enhance inter-view feature alignment. Zero123++ (Shi et al., 2023a) adopts a tiled view representation, enabling single-pass generation over multiple views, a strategy later adopted in Direct2.5 and Instant3D for efficient view synthesis. Similarly, MVDream (Shi et al., 2023b) and Wonder3D (Long et al., 2023) leverage dedicated multi-view self-attention mechanisms to promote cross-view coherence. Meanwhile, other approaches (Chen et al., 2024; Voleti et al., 2024; Yu et al., 2024) exploit spatio-temporal priors from video diffusion models to ensure view consistency across frames.
@@ 2.3 In-Context Learning
Recent advancements in Text-to-Image (T2I) generation (OpenAI, 2023; Podell et al., 2023; Esser et al., 2024; Labs, 2024a) have enabled the synthesis of identity-preserving subject views within a single $M{\times}N$ grid-structured mosaic, where each subview is prompted via carefully designed textual inputs. For example, IC-LoRA (Huang et al., 2024) explores in-context learning by fine-tuning a LoRA (Hu et al., 2021) model on concatenated grid-based image-prompt pairs; however, it suffers from reduced visual consistency, particularly in transferring identity across views. Similarly, by formulating the task as a grid-based image completion problem and replicating the subject image(s) within a structured mosaic layout, (Kang et al., 2025) demonstrates strong identity-preserving capabilities without the need for additional training data, fine-tuning, or modifications during inference.
@@ 3.1 Preliminary
Diffusion Models. Diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020) are a class of likelihood-based generative models that produce data samples via a sequential denoising procedure originating from Gaussian white noise. During training, a predefined forward diffusion process is employed to transform clean observations $\mathbf{x}_{0}$ into a latent noise representation $\mathbf{x}_{T}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ through the iterative injection of Gaussian perturbations across $T$ steps, forming a Markov chain structure, i.e., ${\mathbf{x}}_{t}=\sqrt{\alpha_{t}}{\mathbf{x}}_{0}+\sqrt{1-\alpha_{t}}\epsilon$. The model is trained to learn the backward process, i.e.,
@@
Diffusion Transformers. A growing body of research has begun to adopt transformer (Peebles & Xie, 2023) architectures within text-to-image generative frameworks. Notably, models such as FLUX (Labs, 2024b) exemplify this trend by employing the MultiModal-Diffusion Transformer (MM-DiT) architecture. This design facilitates joint cross-modal interaction by performing attention over concatenated text and image embeddings, thereby enabling more effective integration of multimodal information during the generation process.
@@
Figure 2: Overview of PersonalView. In step 1, we use SMPL (Goel et al., 2023) to fit the body mesh corresponding to the sample from the personalized generator (Guo et al., 2024). Then we render the body mesh for multi-view depth maps. With the in-context depth maps, we can generate the multi-view customization images in step 2 using the personalized model with our control Adapter.
@@ 3.2 Task Formulation
3.3 In-Context Depth-Conditioned Generation
Previous works (Huang et al., 2024; Kang et al., 2025) have demonstrated the contextual learning capabilities of the DiT model, leveraging this to generate images with consistent appearances. Inspired by these advancements, we hypothesize that the DiT model may also possess inherent multi-view geometric consistency. To explore this, we propose utilizing in-context depth maps as cues to activate the multi-view consistency of the DiT model.
Specifically, we adopt a two-stage approach as shown in Figure 2. In the first stage, our objective is to obtain multi-view depth maps that align with the provided prompt. Therefore, we leverage a pre-trained customization generator, like PuLID (Guo et al., 2024), to perform initial sampling based on the user-provided image and prompt. Subsequently, we employ SMPL (Goel et al., 2023) to fit a corresponding human body model and generate multi-view depth maps by rotating and rendering the fitted mesh. In the second stage, these multi-view human depth maps are arranged into a four-panel grid, which serves as an in-context conditioning signal for a pre-trained depth-conditioned model (Labs, 2024b). This setup enables the synthesis of multi-view consistent customized images.
@@ 3.4 Lightweight In-Context Learning
With this in-context depth-conditioned framework, we can further fine-tune the depth-conditioned adapter with an in-context learning paradigm to enhance the geometry and visual consistency. Following FLUX, we adopt LoRA as the adapter mechanism. Empirical results demonstrate that, only a small amount of training samples is required to activate the model’s inherent capability for multi-view consistency. Specifically, we sample multi-view human depth-image pairs from the multi-view dataset like NeRSemble (Kirschstein et al., 2023) and use the VLM model (Bai et al., 2023) to generate captions that correspond to the respective prompts. These depth maps and images are then organized into a four-panel grid format, which is utilized for in-context learning during training. To alleviate the adverse effects of background diversity, we further extract body masks $M$ from the depth maps and integrate them into the diffusion reconstruction loss as spatial priors,
@@
Figure 4: Qualitative comparison. DiffPortrait3D and Era3D exhibit limitations in maintaining geometric and visual consistency, especially with regard to full-body and background regions. Although ViewCrafter achieves improved scene modeling, it does so at the expense of geometric consistency in human representations. Besides, both BAGEL and Qwen-Image demonstrate suboptimal performance in terms of multi-view control. In contrast, our PersonalView achieves superior performance in both geometric fidelity and visual coherence across views.
@@ 3.5 Semantic Correspondence Alignment
While the multi-view consistency is significantly enhanced during the contextual learning process, the relatively limited variety of scenarios in multi-view human data can lead to model overfitting, which in turn diminishes its ability to maintain semantic control. To address this degradation of semantic capabilities, we propose a Semantic Correspondence Alignment Loss as shown in Figure 3.

### v2

# MagicView: Multi-View Consistent Identity Customization via Priors-Guided In-Context Learning

@@ Abstract
Recent advances in personalized generative models have demonstrated impressive capabilities in producing identity-consistent images of the same individual across diverse scenes. However, most existing methods lack explicit viewpoint control and fail to ensure multi-view consistency of generated identities. To address this limitation, we present MagicView, a lightweight adaptation framework that equips existing generative models with multi-view generation capability through 3D priors–guided in-context learning. While prior studies have shown that in-context learning preserves identity consistency across grid samples, its effectiveness in multi-view settings remains unexplored. Building upon this insight, we conduct an in-depth analysis of the multi-view in-context learning ability, and design a conditioning architecture that leverages 3D priors to activate this capability for multi-view consistent identity customization. On the other hand, acquiring robust multi-view capability typically requires large-scale multi-dimensional datasets, which makes incorporating multi-view contextual learning under limited data regimes prone to textual controllability degradation. To address this issue, we introduce a novel Semantic Correspondence Alignment loss, which effectively preserves semantic alignment while maintaining multi-view consistency. Extensive experiments demonstrate that MagicView substantially outperforms recent baselines in multi-view consistency, text alignment, identity similarity, and visual quality, achieving strong results with only 100 multi-view training samples.
Figure 1: MagicView compared to conventional customization methods. MagicView generates personalized images consistent with multiple views given one reference image. Conventional methods like PULID [guo2024pulid] have limited control over the viewpoint in the prompt (i.e., left, middle, and right view) and do not have multi-view consistency.
@@ 1 Introduction
Method Geometry Consistency Visual Consistency w/o Large MV Dataset w/o Test-Time Training Human Identity Consistency Prompt Control
Conventional Customization (e.g., PuLID [guo2024pulid])
Image to Multi-View (e.g., Era3D [li2024era3d])
Unified Multi-Modal Model (e.g., BAGEL [deng2025emerging])
CustomDiffusion360 [kumari2024customizing]
ICLoRA [lhhuang2024iclora]
MagicView
Recent years have seen remarkable progress in human image customization [li2024photomaker, guo2024pulid, wang2024instantid, ipadapter, li2024personalvideo, li2025magicid, yang2024lora], enabling the generation of personalized images from user-provided photographs. Given a single portrait, these methods can synthesize customized depictions of the same individual in diverse contexts, such as sitting on a beach or smiling in a meadow. However, a key challenge remains — achieving multi-view consistent customization that preserves identity and appearance across varying viewpoints.
@@
When users wish to change the view of a customized human while synthesizing it in a novel context (Figure 1), viewpoint-specific prompts such as “from the left view” offer limited control over diverse perspectives. The generated results often lack inter-view coherence, exhibiting inconsistencies in geometry and visual appearance across body shape, facial features, expressions, clothing, and background. This raises a key question: how can multi-view consistency be preserved during personalized image generation?
@@
In this paper, we introduce a new task, Multi-View Customization (MVC), generating multi-view human images conditioned on a single photo. Without test-time tuning, MVC produces identity-faithful results with strong geometry and visual consistency across views, while maintaining flexible prompt-based control (Tab. 1). Beyond personalized creation, this paradigm can be extended to broader applications such as 3D modeling and reconstruction.
@@
The core of this task is how to control the generation of multiple views while maintaining geometric and visual consistency. We are inspired by the in-context learning ability of the DiT-based models, such as FLUX [flux1-dev]. Prior works [lhhuang2024iclora, kang2025flux] have demonstrated the ability to generate a grid of images with roughly consistent visual content. However, they unexplore the potential to preserve the consistency in multi-view scenarios. It motivates us to investigate whether the model has a strong multi-view capability to foster more explicit multi-view consistency.
@@
To this end, we employ in-context depth maps as 3D priors to activate FLUX’s capability for geometrically and visually consistent generation. Specifically, we propose a multi-view depth–based framework, termed MagicView. Our goal is to generate multi-view depth maps consistent with the given prompts, serving as contextual priors for subsequent synthesis. In the first stage, we leverage a pre-trained customization model [guo2024pulid] for preliminary sampling and use the SMPL model [goel2023humans] to fit and render multi-view depth maps. These depth maps are then arranged into a four-panel grid, providing in-context conditioning for a depth-guided model to synthesize multi-view consistent customized images.
However, acquiring robust multi-view customization capability typically demands large-scale multi-dimensional datasets, making multi-view contextual learning under limited data prone to textual controllability degradation. To address this issue, we propose a Semantic Correspondence Alignment (SCA) loss tailored for DiT-based models to preserve the model’s intrinsic semantic controllability. Specifically, SCA incorporates a frozen reference branch of the original model during training and enforces token-level correspondence alignment between the trainable and frozen branches across both text and visual modalities. This design effectively maintains the model’s semantic precision while preserving its in-context learning capability for multi-view reasoning.
@@
Our proposed MagicView consistently preserves multi-view human identity requiring only a few training steps without large-scale multi-view human dataset. It exhibits robust multi-view consistency across various attributes, including facial features, body gestures, background, and clothing, while concurrently preserving excellent prompt controllability and identity fidelity. Extensive experimentation demonstrates its superior performance compared to existing novel view synthesis methods that typically necessitate extensive training on large-scale multi-view datasets.
@@
  • We introduce a novel task, Multi-View Customization conditioned on a single user-provided photograph without test-time tuning, which aims to generate multi-view consistent customized images with precise identity and diverse prompt-control.
  • We propose MagicView, leveraging in-context multiview priors to activate the in-context learning capabilities of DiT-based models with only a few training samples, facilitating the generation of geometry and visual consistent human images.
  • We introduce a Semantic Correspondence Alignment loss tailored for DiT-based models, which preserves the original model’s semantic control capabilities with limited multi-view data.
@@ 2.1 Image Customization
In the domain of Text-to-Image (T2I) generation, a variety of approaches have emerged to address identity (ID) customization [gal2022image, li2024photomaker, gal2024lcm, valevski2023face0, xiao2024fastcomposer, ma2024subject, peng2024portraitbooth, li2023few, li2024unihda, chen2023photoverse, jin2025scar, xu2025scalar, lan2025flux]. A seminal method in this line of work is Textual Inversion [gal2022image], which encodes user-specific identity information into a dedicated token embedding while keeping the T2I model parameters fixed. To improve identity fidelity, In contrast, encoder-based paradigms aim to directly inject identity representations into the generation pipeline. For instance, PhotoMaker [li2024photomaker] leverages large-scale identity datasets to construct robust ID embeddings from diverse image samples. Similarly, PuLID [guo2024pulid] introduces a more precise ID supervision mechanism by minimizing identity loss between synthesized outputs and reference images. Recently, CustomDiffusion360 enables explicit control over object viewpoints in the customization of text-to-image diffusion models; however, it requires test-time training and is limited to object-level customization rather than human subjects.
@@ 2.2 Multi-View Images Synthesis
Cross-view consistency plays a pivotal role in multi-view generation. MVFusion [tang2023MVDiffusion] initiates this direction by parallel multi-view image generation with correspondence-aware attention, which facilitates cross-view information exchange and supports textured scene mesh reconstruction. Building upon it, subsequent works [tseng2023consistent, kant2024spad, gu2024diffportrait3d, li2024gca, shen2024imagpose] incorporate epipolar constraints into diffusion models to enhance inter-view feature alignment. Zero123++ [shi2023zero123++] adopts a tiled view representation, enabling single-pass generation over multiple views, a strategy later adopted in Direct2.5 and Instant3D for efficient view synthesis. Similarly, MVDream [shi2023mvdream] and Wonder3D [long2023wonder3d] leverage dedicated multi-view self-attention mechanisms to promote cross-view coherence. Meanwhile, other approaches [chen2024v3d, voleti2024sv3d, yu2024viewcrafter] exploit spatio-temporal priors from video diffusion models to ensure view consistency across frames.
@@ 2.3 In-Context Learning
Recent advancements in Text-to-Image (T2I) generation [openai2023dalle3, podell2023sdxl, esser2024scaling, black2024flux] have enabled the synthesis of identity-preserving subject views within a single $M{\times}N$ grid-structured mosaic, where each subview is prompted via carefully designed textual inputs. For example, by formulating the task as a grid-based image completion problem and replicating the subject image(s) within a structured mosaic layout, [kang2025flux] demonstrates strong identity-preserving capabilities without the need for additional training data, fine-tuning, or modifications during inference. IC-LoRA [lhhuang2024iclora] explores in-context learning by fine-tuning a LoRA [hu2021lora] model on concatenated grid-based image-prompt pairs. However, they under-explore the multi-view scenarios which suffers from reduced visual consistency in transferring identity across views.
@@ 3 Method
Figure 2: Overview of MagicView. In step 1, we use SMPL [goel2023humans] to fit the body mesh corresponding to the sample from the personalized generator [guo2024pulid]. Then we render the body mesh for multi-view depth maps. With the in-context depth priors, we can generate the multi-view customization images in step 2 using the personalized model with our control Adapter.
@@ 3.1 Preliminary
Diffusion Models. Diffusion models [sohl2015deep, ho2020denoising] are a class of likelihood-based generative models that produce data samples via a sequential denoising procedure originating from Gaussian white noise. During training, a predefined forward diffusion process is employed to transform clean observations $\mathbf{x}_{0}$ into a latent noise representation $\mathbf{x}_{T}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ through the iterative injection of Gaussian perturbations across $T$ steps, forming a Markov chain structure, i.e., ${\mathbf{x}}_{t}=\sqrt{\alpha_{t}}{\mathbf{x}}_{0}+\sqrt{1-\alpha_{t}}\epsilon$. The model is trained to learn the backward process, i.e.,
@@
Diffusion Transformers. A growing body of research has begun to adopt transformer [dit] architectures within text-to-image generative frameworks. Notably, models such as FLUX [flux1-dev] exemplify this trend by employing the MultiModal-Diffusion Transformer (MM-DiT) architecture. This design facilitates joint cross-modal interaction by performing attention over concatenated text and image embeddings, thereby enabling more effective integration of multimodal information during the generation process.
@@ 3.2 Task Formulation
3.3 Priors-Guided In-Context Generation
Previous works [lhhuang2024iclora, kang2025flux] have demonstrated the contextual learning capabilities of the DiT model, leveraging this to generate images with consistent appearances. However, they unexplore the potential to preserve the consistency in multi-view scenarios. It motivates us to investigate whether the model has a strong multi-view capability to foster more explicit multi-view consistency. To explore this, we propose utilizing in-context depth maps as priors to activate the multi-view consistency of the DiT model.
Specifically, we adopt a two-stage approach as shown in Figure 2. In the first stage, our objective is to obtain multi-view depth maps that align with the provided prompt. Therefore, we leverage a pre-trained customization generator, like PuLID [guo2024pulid], to perform initial sampling based on the user-provided image and prompt. Subsequently, we employ SMPL [goel2023humans] to fit a corresponding human body model and generate multi-view depth maps by rotating and rendering the fitted mesh. In the second stage, these multi-view human depth maps are arranged into a four-panel grid, which serves as an in-context conditioning signal for a pre-trained depth-conditioned model [flux1-dev]. This setup enables the synthesis of multi-view consistent customized images.
@@ 3.4 Lightweight In-Context Learning
With this in-context depth-conditioned framework, we can further fine-tune the depth-conditioned adapter with an in-context learning paradigm to enhance the geometry and visual consistency. Empirical results demonstrate that, only a small amount of training samples is required to activate the model’s inherent capability for multi-view consistency. Specifically, we sample multi-view human depth-image pairs from the multi-view dataset like NeRSemble [kirschstein2023nersemble] and use the VLM model [bai2023qwenvlversatilevisionlanguagemodel] to generate captions that correspond to the respective prompts. These depth maps and images are then organized into a four-panel grid format, which is utilized for in-context learning during training. To alleviate the adverse effects of background diversity, we further extract body masks $M$ from the depth maps and integrate them into the diffusion reconstruction loss as spatial priors,
@@ 3.5 Semantic Correspondence Alignment
While the multi-view consistency is significantly activated with the 3D priors, acquiring robust multi-view customization capability typically demands large-scale multi-dimensional datasets, making multi-view contextual learning under limited data prone to textual controllability degradation. To address this degradation of semantic capabilities, we propose a Semantic Correspondence Alignment Loss as shown in Figure 3.
@@
Figure 4: Qualitative comparison. DiffPortrait3D and Era3D exhibit limitations in maintaining geometric and visual consistency, especially with regard to full-body and background regions. Although ViewCrafter achieves improved scene modeling, it does so at the expense of geometric consistency in human representations. Besides, both BAGEL and Qwen-Image demonstrate suboptimal performance in terms of multi-view control. In contrast, our MagicView achieves superior performance in both geometric fidelity and visual coherence across views. The baseline methods generate results using prompts that explicitly specify viewpoints, such as “left view”, “middle view”, and “right view”.